OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 2x - y = -8 3x + 2y = 2

OpenStudy (anonymous):

y=2x+8(from the 1st one) put it in the second one: 3x+2(2x+8)=2 7x=-14 x=-2 so, y=2x+8=2(-2)+8=4